OpenStudy (samigupta8):

The expression [(1+sinπ/8 + icosπ/8)/(1+sinπ/8 -icosπ/8)]^8 =?

OpenStudy (vishweshshrimali5):

You can use Euler's formula for representing complex numbers... \[e^{i\theta} = \cos{\theta} + i \sin{\theta}\] Also, use the fact that, \(\cos{\pi/2 - \theta} = \sin{\theta}\)

OpenStudy (anonymous):

Is it \[\cos(\frac{ \pi }{ 8 }) or \frac{ \cos (\pi) }{ 8 }\]

OpenStudy (vishweshshrimali5):

Sorry that is -> \(\cos(\pi/2 - \theta) = \sin{\theta}\) and \(\sin(\pi/2 - \theta) = \cos{\theta}\)

OpenStudy (samigupta8):

Former one ...

OpenStudy (vishweshshrimali5):

Can you convert it into Euler formula?

OpenStudy (samigupta8):

I used this 1+i (cos theta-isin theta)/1-i(cos theta+isintheta)

OpenStudy (samigupta8):

That is 1+ie^-itheta/1-ie^itheta

OpenStudy (vishweshshrimali5):

\[\cfrac{1 + ie^{-i \theta}}{1 - ie^{i \theta}}\] Multiply numerator and denominator by \(e^{i\theta}\)

OpenStudy (vishweshshrimali5):

Wait don't multiply it :P Not gonna help much...

OpenStudy (samigupta8):

I got this i(e^itheta +e^-itheta)/(1+e^itheta)

OpenStudy (samigupta8):

By rationalisation

OpenStudy (samigupta8):

Actually no..a slight mistake

OpenStudy (samigupta8):

It wud be 1+e^i2theta in denominator

OpenStudy (samigupta8):

I came up with this as ie^-itheta

OpenStudy (vishweshshrimali5):

See we need to convert it to pure e^i theta type term... Then only we would be able to apply power 8... Now if only I can think of a way of doing that.. :/

OpenStudy (samigupta8):

Got it.....thanks..

OpenStudy (vishweshshrimali5):

@hartnn any ideas?

OpenStudy (samigupta8):

My ans is -1..

OpenStudy (vishweshshrimali5):

Perfect!!

OpenStudy (vishweshshrimali5):

:P Care to share how you get it ?

OpenStudy (vishweshshrimali5):

*got

OpenStudy (samigupta8):

Okay.. Let's begin then!

OpenStudy (samigupta8):

I got till that point i(e^itheta+e^-itheta)/(1+e^i2theta) Now if we simplify this numerator term is nothing but 2cos theta And denominator is 2costheta(costheta+isintheta) So we r left with this 1/e^itheta That is e^-itheta Now there is a power of 8 on it.. So we get e^-i8theta Plug in the value of theta as π/8 We get the ans as -1